OpenStudy (anonymous):

A square pyramid has a slant height of 5 feet and a height of 4 feet. What is the surface area if the square base has sides of length 6 feet? Can you explain it? Thanks!

OpenStudy (anonymous):

Find the area of the base first: \[Base(Area) = Side^2\] Side is given as 6.. Can you find the base area.???

OpenStudy (anonymous):

So it would be Base(A) = 6^2 making 36?

OpenStudy (anonymous):

Yes you are right..

OpenStudy (anonymous):

What next?

OpenStudy (anonymous):

Now Surface Area of Square Pyramid is given by: \[\large \color{green}{SA = Base \;Area + 2 \times s \times l}\]

OpenStudy (anonymous):

Here you have found the base area.. s is side given that is 6.. l is the slant height that is 5.. Can you plug in the values to solve..

OpenStudy (anonymous):

Alright gimie a sec.

OpenStudy (anonymous):

Remember one thing: We are given with Height too that is 4 feet but to find the surface area we need to haver height.. Yeah that height is surely used in finding the volume of Square Pyramid..

OpenStudy (anonymous):

Take your time..

OpenStudy (anonymous):

SA= 96?

OpenStudy (anonymous):

Absolutely.. Well done..

OpenStudy (anonymous):

Don't forget to use the units that for Area are \(feet^2\)..

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

Did you understand the whole thing??

OpenStudy (anonymous):

I want you too remember the formula for Surface Area so that next time you can do it yourself only..

OpenStudy (anonymous):

*to in place of too

OpenStudy (anonymous):

Good Luck.. Welcome dear..

OpenStudy (anonymous):

Yeah you find the BA the plug the values in and it was 96FT^2

OpenStudy (anonymous):

Yes... Absolutely.. You are a good learner too..